The sum of the numbers on two adjacent post office boxes is 697, What are the numbers?
We are suppose to write an equation and solve.
I just dont know how to figure this out.

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

The sum of the numbers on two adjacent post office boxes is 697, What are the numbers?
.
Let x = first box number
then because the second is adjacent:
x+1 = adjacent box number
.
x + x+1 = 697
2x+1 = 697
2x = 696
x = 348
.
adjacent box:
x+1 = 348+1 = 349
.
answer: 349 and 348

Answer by Alan3354(69443)

697/2 = 348.5, the average, or center
--> 348 & 349